# DET-LSH: A Locality-Sensitive Hashing Scheme with Dynamic Encoding Tree for Approximate Nearest Neighbor Search

### v1

@@ 1. Introduction
Limitations and Motivation. Efficiency and accuracy are key metrics to evaluate the performance of LSH-based methods in $c$-ANN search. Nowadays, new data is produced at an ever-increasing rate, and the size of datasets is continuously growing (Palpanas, 2015; Palpanas and Beckmann, 2019; Fernandez et al., 2020; Wei et al., 2023). We need to manage large-scale data more efficiently to support further data analysis (Wellenzohn et al., 2023; Peng et al., 2023, 2022; Echihabi et al., 2019). However, existing LSH-based methods mainly focus on reducing query time and improving query accuracy by designing different query strategies, but pay little attention to reducing indexing time (Tian et al., 2023b; Huang et al., 2015; Lu and Kudo, 2020; Lu et al., 2020; Sun et al., 2014; Zheng et al., 2020). They typically fine-tune existing data-oriented partitioning trees to index data points and support their query strategies, such as R*-Tree (Beckmann et al., 1990) for DB-LSH (Tian et al., 2023b), PM-Tree (Skopal et al., 2005) for PM-LSH (Zheng et al., 2020), and R-Tree (Guttman, 1984) for SRS (Sun et al., 2014). Data-oriented partitioning trees (Guttman, 1984; Ciaccia et al., 1997; Beckmann et al., 1990; Skopal et al., 2005) group nearby data points and partition them into their minimum bounding objects (e.g., hyperrectangle, hypersphere) hierarchically. However, partitioning directly in a multi-dimensional space is time-consuming, which limits the efficiency of these methods in the indexing phase. In addition, the performance of data-oriented partitioning trees decreases as the space dimensionality increases (Böhm, 2000; Weber et al., 1998), which limits the dimensionality of the projected space. Therefore, it is necessary to design a more efficient tree structure to address the limitations. From another perspective, a more efficient tree structure can also help improve query accuracy, since more trees can be constructed in the same indexing time, and query answering based on more trees can be more accurate. For example, the state-of-the-art method among BC methods, DB-LSH (Tian et al., 2023b), constructs five R*-Trees to reduce the probability of missing exact NN points in the query phase.
@@ 3.1. Problem Definition
$o^{\prime},q^{\prime}$ $o$ and $q$ in the projected space
@@
$\mathcal{H}(o)$ $[h_{1}(o),...,h_{K}(o)]$, the coordinates of $o^{\prime}$
@@ 3.2. Locality-Sensitive Hashing
Given a distance $r$, an approximation ratio $c>1$, a family of hash functions =$\mathcal{H}=\{h:\mathbb{R}^{d}\rightarrow\mathbb{R}\}$ is called ($r$,$cr$,$p_{1}$,$p_{2}$)-locality-sensitive, if for $\forall o_{1},o_{2}\in\mathbb{R}^{d}$, it satisfies both of the following conditions:
@@
3.3. $p$-Stable Distribution and $\chi^{2}$ Distribution
@@ 4. The DET-LSH Method
In this section, we present the details of DET-LSH and the design of Dynamic Encoding Tree (DE-Tree). DET-LSH consists of three phases: an encoding phase to encode the LSH-based projected points into iSAX representations; an indexing phase to construct DE-Trees based on the iSAX representations; a query phase to perform range queries in DE-Trees for ANN search. Figure LABEL:overview provides a high-level overview of the workflow for DET-LSH.
@@ 4.1. Encoding Phase
Static encoding scheme. In data series similarity search, traditional iSAX-based methods adopt the static encoding scheme (Camerra et al., 2014; Zoumpatianos et al., 2016; Peng et al., 2020a, b, 2021b; Chatzakis et al., 2023; Fatourou et al., 2023). Since normalized data series have highly Gaussian distribution (Shieh and Keogh, 2008), they simply determine the breakpoints $b_{1},...,b_{a-1}$ such that the area under a $\mathcal{N}(0,1)$ Gaussian curve from $b_{i}$ to $b_{i+1}$ is $\frac{1}{a}$, where $b_{0}$ and $b_{a}$ are defined as $\infty$- ∞ and +$\infty$+ ∞. Therefore, these breakpoints are static and independent of datasets. Existing methods encode a data series by checking which two breakpoints each of its coordinates falls between in a common statistical table. However, the datasets for ANN search have arbitrary distributions, so the static encoding scheme is no longer suitable.
Dynamic encoding scheme. In DET-LSH, we design a dynamic encoding scheme to dynamically select breakpoints based on the distribution of the dataset, aiming to divide data points into different regions as evenly as possible, i.e., each region contains the same number of points. Specifically, assuming we have a dataset with cardinality $n$, we first use $K\cdot L$ hash functions to calculate the $K$-dimensional points in $L$ projected spaces, where $\mathcal{H}_{i}(o)=[h_{i1}(o),...,h_{iK}(o)]$ denote a point $o$ in the $i$-th projected space. Let $C_{ij}=[h_{ij}(o_{1}),...,h_{ij}(o_{n})]$ denote the set of coordinates of all $n$ points in their $i$-th projected space and $j$-th dimension, where $i=1,...,L$ and $j=1,...,K$. We denote $C^{\uparrow}_{ij}$ as a new set in which the elements of $C_{ij}$ are sorted in ascending order, and use $C^{\uparrow}_{ij}(t)$ to represent the $t$-th element in $C^{\uparrow}_{ij}$. Intuitively, to make points evenly divided into $N_{r}=256$ regions in each dimension, we can select ordered breakpoints $B_{ij}$ from $C^{\uparrow}_{ij}$, where $B_{ij}(z)=C^{\uparrow}_{ij}(\lfloor\frac{n}{N_{r}}\rfloor\cdot(z-1))$ and $z=2,...,N_{r}$. We set $B_{ij}(1)=C^{\uparrow}_{ij}(1)$ and $B_{ij}(N_{r}+1)=C^{\uparrow}_{ij}(n)$. In practice, we dynamically select $N_{r}+1$ breakpoints for each dimension. Then, for any point $o$, each dimension $h_{ij}(o)$ can be independently encoded based on the selected breakpoints $B_{ij}$.
@@
Output: A set of encoded points $EP$
1Initialize $EP$ with size $n\cdot L\cdot K$;
2 $B\leftarrow$ call BreakpointsSelection($K$,$L$,$n$,$P$,$n_{s}$,$N_{r}$);
@@ 8                   $EP_{ij}(o_{z})\leftarrow b$-th symbol in the 8-bit alphabet;
In terms of algorithm design, the intuitive idea is to completely sort $C_{ij}$ to get the exact $C^{\uparrow}_{ij}$ and then select breakpoints from it. However, we only need $N_{r}+1$ discrete elements in $C^{\uparrow}_{ij}$, the complete sorting of $C_{ij}$ is wasteful. Therefore, combining the QuickSelect algorithm with the divide-and-conquer strategy, we design a dynamic encoding scheme based on the unordered $C_{ij}$. Algorithm 1 introduces how we dynamically select breakpoints, which is the first step of the encoding scheme. To improve efficiency, we randomly sample $n_{s}$ points from the dataset and select breakpoints based on these sampled points. In practice, we set $n_{s}=0.1n$. For each $C_{ij}$, Algorithm 1 obtains breakpoints by running multiple rounds of the QuickSelect algorithm combined with the divide-and-conquer strategy (lines 6-8). For unordered $C_{ij}$, QuickSelect($start$, $q$, $end$) can find the $q$-th smallest element between $C_{ij}(start)$ and $C_{ij}(end)$, and move it to the position of $C_{ij}(start+q)$. Then, $C_{ij}(start+q)$ is greater than all elements from $C_{ij}(start)$ to $C_{ij}(start+q-1)$ and smaller than all elements from $C_{ij}(start+q+1)$ to $C_{ij}(end)$. Therefore, we can select a single breakpoint from $C_{ij}$ by running QuickSelect once. Since we set $N_{r}=256$, the divide-and-conquer strategy can be perfectly applied to our algorithm. Specifically, a total of $\log_{2}N_{r}$ rounds need to run, and the $z$-th round select $2^{z-1}$ breakpoints by running QuickSelect in $2^{z-1}$ sub-regions generated from the ($z$-1)-th round, where $z=1,...,\log_{2}N_{r}$. For each $C_{ij}$, we select the minimum element as the first breakpoint $B_{ij}(1)$ and the maximum element as the last breakpoint $B_{ij}(N_{r}+1)$ (lines 9-11). In practice, Algorithm 1 achieves 3x speedup in running time over the complete sorting scheme, as shown in Section 6.2. After getting the breakpoint set $B$, Algorithm 2 will encode all points into iSAX representations and return the set of encoded points $EP$ (lines 3-8).
Input: Parameters $K$, $L$, $n$, encoded points set $EP$, maximum size of a leaf node $max\_size$
@@ 6             $target\_leaf\leftarrow$ leaf node of $T_{i}$ to insert $\langle ep_{i}(o_{z}),pos_{z}\rangle$;
8                   SplitNode($target\_leaf$);
@@ 9                   $target\_leaf\leftarrow$ the new leaf node to insert $\langle ep_{i}(o_{z}),pos_{z}\rangle$;
11            Insert $\langle ep_{i}(o_{z}),pos_{z}\rangle$ to $target\_leaf$;
@@ 3       $node\leftarrow$ the $i$-th child of root node in $T$;
4       call TraverseSubtree($node$, $q^{\prime}$, $r^{\prime}$, $S$);
@@
Input: A node $node$, the projected query point $q^{\prime}$, the search radius $r^{\prime}$, the set of points $S$
@@
5else if $node$ is a leaf then
6       $upper\_bound\_dist\leftarrow$ calculate the upper bound distance between $q^{\prime}$ and $node$;
@@
8             $S\leftarrow S\,\cup$ all points in $node$;
@@
11             while $node$ has next point do
@@ 15                         $S\leftarrow S\,\cup o$;
In a DE-Tree, except the root node that has $2^{K}$ children, other internal nodes have only two children. This is because when an internal node needs to be split, we only select one of its $K$ dimensions for further bit-wise binary division. For example, in Figure 1(b), we choose the first dimension of node [$0^{*}$,$0^{*}$] to split, and the representations of its two children are [$0^{*}$] and [$0^{*}$]. The choice of which dimension to divide is important for splitting nodes. Intuitively, splitting a node works better if the obtained two children contain similar numbers of points. Therefore, when splitting nodes, we choose the dimension that most evenly divides the points.
@@ 5       $S\leftarrow S\cup S_{i}$;
10if |$\lvert\left\{o\mid o\in S\land\left\|o,q\right\|\leq c\cdot r\right\}\rvert\geq 1$ then
@@ 6             $S\leftarrow S\cup S_{i}$;
11      if |$\lvert\left\{o\mid o\in S\land\left\|o,q\right\|\leq c\cdot r\right\}\rvert\geq k$ then
@@ 5.1. Quality Guarantee
Let $\mathcal{H}_{i}(o)=[h_{i1}(o),...,h_{iK}(o)]$ denote a data point $o$ in the $i$-th projected space, where $i=1,...,L$. We define three events as follows:
  • E1: If there exists a point $o$ satisfying $\left\|o,q\right\|\leq r$, then its projected distance to $q$, i.e., $\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$, is smaller than $\epsilon r$ for some $i=1,...,L$;
  • E2: If there exists a point $o$ satisfying $\left\|o,q\right\|>cr$, then its projected distance to $q$, i.e., $\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$, is smaller than $\epsilon r$ for some $i=1,...,L$;
@@
Given a point $o$ satisfying $\left\|o,q\right\|\leq r$, let $s=\left\|o,q\right\|$ and $s_{i}^{\prime}=\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$ denote the distances between $o$ and $q$ in the original space and in the $i$-th projected space, where $i=1,...,L$. From Equation 3, we have $\sqrt{\chi^{2}_{\alpha_{1}}(K)}=\epsilon$. For each independent projected space, from Lemma 6, we have $\Pr{[s_{i}^{\prime}>s\sqrt{\chi^{2}_{\alpha_{1}}(K)}]}=\Pr{[s_{i}^{\prime}>\epsilon s]}=\alpha_{1}$. Since $s\leq r$, $\Pr{[s_{i}^{\prime}>\epsilon r]}\leq\alpha_{1}$. Considering $L$ projected spaces, we have $\Pr{[\textbf{E1}]}\geq 1-\alpha_{1}^{L}=1-\frac{1}{\mathrm{e}}$. Likewise, given a point $o$ satisfying $\left\|o,q\right\|>cr$, let $s=\left\|o,q\right\|$ and $s_{i}^{\prime}=\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$ denote the distances between $o$ and $q$ in the original space and in the $i$-th projected space, where $i=1,...,L$. From Equation 3, we have $\sqrt{\chi^{2}_{\alpha_{2}}(K)}=\frac{\epsilon}{c}$. For each independent projected space, from Lemma 6, we have $\Pr{[s_{i}^{\prime}>s\sqrt{\chi^{2}_{\alpha_{2}}(K)}]}=\Pr{[s_{i}^{\prime}>\frac{\epsilon s}{c}]}=\alpha_{2}$. Since $s>cr$, i.e., $\frac{s}{c}>r$, $\Pr{[s_{i}^{\prime}>\epsilon r]}>\alpha_{2}$. Considering $L$ projected spaces, we have $\Pr{[\textbf{E2}]}\leq 1-\alpha_{2}^{L}$, thus the expected number of such points in dataset $\mathcal{D}$ is upper bounded by $(1-\alpha_{2}^{L})\cdot n$. By Markov’s inequality, we have $\Pr{[\textbf{E3}]}>1-\frac{(1-\alpha_{2}^{L})\cdot n}{\beta n}=\frac{1}{2}.$
@@
We show that when E1 and E3 hold at the same time, Algorithm 7 returns a correct $c^{2}$-$k$-ANN result. Let $o_{i}^{*}$ be the $i$-th exact nearest point to $q$ in $\mathcal{D}$, we assume that $r_{i}^{*}=\left\|o_{i}^{*},q\right\|>r_{min}$, where $r_{min}$ is the initial search radius and $i=1,...,k$. We denote the number of points in the candidate set under search radius $r$ as $\lvert S_{r}\rvert$. Obviously, when enlarging the search radius $r=r_{min},r_{min}\cdot c,r_{min}\cdot c^{2},...$, there must exist a radius $r_{0}$ satisfying $\lvert S_{r_{0}}\rvert<\beta n+k$ and $\lvert S_{c\cdot r_{0}}\rvert\geq\beta n+k$. The distribution of $r_{i}^{*}$ has three cases:
@@
    Case 1: If for all $i=1,...,k$ satisfying $r_{i}^{*}\leq r_{0}$, which indicates the range queries in all $L$ index trees have been executed at $r=r_{0}$ (lines 3-8). Due to E1, all $r_{i}^{*}$ must in $S_{r_{0}}$. Since $S_{r_{0}}\subsetneqq S_{c\cdot r_{0}}$, all $r_{i}^{*}$ also must in $S_{c\cdot r_{0}}$. Therefore, Algorithm 7 returns the exact $k$ nearest points $o_{i}^{*}$ to $q$.
@@
    Case 2: If for all $i=1,...,k$ satisfying $r_{i}^{*}>r_{0}$, all $r_{i}^{*}$ not belong to $S_{r_{0}}$. Since Algorithm 7 may terminate after executing range queries in part of $L$ index trees at $r=c\cdot r_{0}$ (line 8), we cannot guarantee that $r_{i}^{*}\leq c\cdot r_{0}$. However, due to E3, there are at least $k$ points $o_{i}$ in $S_{c\cdot r_{0}}$ satisfying $\left\|o_{i},q\right\|\leq c^{2}r_{0}$, $i=1,...,k$. Therefore, we have $\left\|o_{i},q\right\|\leq c^{2}r_{0}\leq c^{2}r_{i}^{*}$, i.e., each $o_{i}$ is a $c^{2}$-ANN point for corresponding $o_{i}^{*}$.
@@
    Case 3: If there exists an integer $m\in(1,k)$ such that for all $i=1,...,m$ satisfying $r_{i}^{*}\leq r_{0}$ and for all $i=m+1,...,k$ satisfying $r_{i}^{*}>r_{0}$, indicating that Case 3 is a combination of Case 1 and Case 2. For each $i\in[1,m]$, Algorithm 7 returns the exact nearest point $o_{i}^{*}$ to $q$ based on Case 1. For each $i\in[m+1,k]$, Algorithm 7 returns a $c^{2}$-ANN point for $o_{i}^{*}$ based on Case 2.

### v2

@@ 1. Introduction
Limitations and Motivation. Efficiency and accuracy are key metrics to evaluate the performance of LSH-based methods in $c$-ANN search. Nowadays, new data is produced at an ever-increasing rate, and the size of datasets is continuously growing (Palpanas, 2015; Palpanas and Beckmann, 48(3), 2019; Fernandez et al., 2020; Wei et al., 2023). We need to manage large-scale data more efficiently to support further data analysis (Wellenzohn et al., 2023; Peng et al., 2023, 2022; Echihabi et al., 2019). However, existing LSH-based methods mainly focus on reducing query time and improving query accuracy by designing different query strategies, but pay little attention to reducing indexing time (Tian et al., 2023b; Huang et al., 2015; Lu and Kudo, 2020; Lu et al., 2020; Sun et al., 2014; Zheng et al., 2020). They typically fine-tune existing data-oriented partitioning trees to index data points and support their query strategies, such as R*-Tree (Beckmann et al., 1990) for DB-LSH (Tian et al., 2023b), PM-Tree (Skopal et al., 2005) for PM-LSH (Zheng et al., 2020), and R-Tree (Guttman, 1984) for SRS (Sun et al., 2014). Data-oriented partitioning trees (Guttman, 1984; Ciaccia et al., 1997; Beckmann et al., 1990; Skopal et al., 2005) group nearby data points and partition them into their minimum bounding objects (e.g., hyperrectangle, hypersphere) hierarchically. However, partitioning directly in a multi-dimensional space is time-consuming, which limits the efficiency of these methods in the indexing phase. In addition, the performance of data-oriented partitioning trees decreases as the space dimensionality increases (Böhm, 2000; Weber et al., 1998), which limits the dimensionality of the projected space. Therefore, it is necessary to design a more efficient tree structure to address the limitations. From another perspective, a more efficient tree structure can also help improve query accuracy, since more trees can be constructed in the same indexing time, and query answering based on more trees can be more accurate. For example, the state-of-the-art method among BC methods, DB-LSH (Tian et al., 2023b), constructs five R*-Trees to reduce the probability of missing exact NN points in the query phase.
@@ 3.1. Problem Definition
$^{\prime},q^{\prime} oo$ and $q$ in the projected space
@@
$\mathcal{H}(o)$ $h_{1}(o),...,h_{K}(o)]$, the coordinates of $o^{\prime}$
@@
(a) Encoding phase and indexing phase.
(b) Query phase.
Figure 2. Overview of the DET-LSH workflow.
@@ 3.2. Locality-Sensitive Hashing
[…]
@@
$3.$ $p$-Stable Distribution and $\chi^{2}$ Distribution
@@ 4. The DET-LSH Method
In this section, we present the details of DET-LSH and the design of Dynamic Encoding Tree (DE-Tree). DET-LSH consists of three phases: an encoding phase to encode the LSH-based projected points into iSAX representations; an indexing phase to construct DE-Trees based on the iSAX representations; a query phase to perform range queries in DE-Trees for ANN search. Figure 2 provides a high-level overview of the workflow for DET-LSH.
@@ 4.1. Encoding Phase
Static encoding scheme. In data series similarity search, traditional iSAX-based methods adopt the static encoding scheme (Camerra et al., 2014; Zoumpatianos et al., 2016; Peng et al., 2020a, b, 2021b; Chatzakis et al., 2023; Fatourou et al., 2023). Since normalized data series have highly Gaussian distribution (Shieh and Keogh, 2008), they simply determine the breakpoints $b_{1},...,b_{a-1}$ such that the area under a $\mathcal{N}(0,1)$ Gaussian curve from $b_{i}$ to $b_{i+1}$ is $\frac{1}{a}$, where $b_{0}$ and $b_{a}$ are defined as -$\infty$ and $\infty$. Therefore, these breakpoints are static and independent of datasets. Existing methods encode a data series by checking which two breakpoints each of its coordinates falls between in a common statistical table. However, the datasets for ANN search have arbitrary distributions, so the static encoding scheme is no longer suitable.
Dynamic encoding scheme. In DET-LSH, we design a dynamic encoding scheme to dynamically select breakpoints based on the distribution of the dataset, aiming to divide data points into different regions as evenly as possible, i.e., each region contains the same number of points. Specifically, assuming we have a dataset with cardinality $n$, we first use $K\cdot L$ hash functions to calculate the $K$-dimensional points in $L$ projected spaces, where $\mathcal{H}_{i}(o)=[h_{i1}(o),...,h_{iK}(o)]$ denote a point $o$ in the $i$-th projected space. Let $C_{ij}=[h_{ij}(o_{1}),...,h_{ij}(o_{n})]$ denote the set of coordinates of all $n$ points in their $i$-th projected space and $j$-th dimension, where i=1,,Li=1,...,L and j=1,,Kj=1,...,K. We denote $C^{\uparrow}_{ij}$ as a new set in which the elements of $C_{ij}$ are sorted in ascending order, and use $C^{\uparrow}_{ij}(t)$ to represent the $t$-th element in $C^{\uparrow}_{ij}$. Intuitively, to make points evenly divided into $N_{r}=256$ regions in each dimension, we can select ordered breakpoints $B_{ij}$ from $C^{\uparrow}_{ij}$, where $B_{ij}(z)=C^{\uparrow}_{ij}(\lfloor\frac{n}{N_{r}}\rfloor\cdot(z-1))$ and $z=2,...,N_{r}$. We set $B_{ij}(1)=C^{\uparrow}_{ij}(1)$ and $B_{ij}(N_{r}+1)=C^{\uparrow}_{ij}(n)$. In practice, we dynamically select $N_{r}+1$ breakpoints for each dimension. Then, for any point $o$, each dimension $h_{ij}(o)$ can be independently encoded based on the selected breakpoints $B_{ij}$.
@@
Output: A set of encoded points EPEP
1Initialize EPEP with size $n\cdot L\cdot K$;
[…]
@@ 8       $EP_{ij}(o_{z})\leftarrow b$-th symbol in the 8-bit alphabet;
In terms of algorithm design, the intuitive idea is to completely sort $C_{ij}$ to get the exact $C^{\uparrow}_{ij}$ and then select breakpoints from it. However, we only need $N_{r}+1$ discrete elements in $C^{\uparrow}_{ij}$, the complete sorting of $C_{ij}$ is wasteful. Therefore, combining the QuickSelect algorithm with the divide-and-conquer strategy, we design a dynamic encoding scheme based on the unordered $C_{ij}$. Algorithm 1 introduces how we dynamically select breakpoints, which is the first step of the encoding scheme. To improve efficiency, we randomly sample $n_{s}$ points from the dataset and select breakpoints based on these sampled points. In practice, we set $n_{s}=0.1n$. For each $C_{ij}$, Algorithm 1 obtains breakpoints by running multiple rounds of the QuickSelect algorithm combined with the divide-and-conquer strategy (lines 6-8). For unordered $C_{ij}$, QuickSelect(startstart, $q$, endend) can find the $q$-th smallest element between $C_{ij}(start)$ and $C_{ij}(end)$, and move it to the position of $C_{ij}(start+q)$. Then, $C_{ij}(start+q)$ is greater than all elements from $C_{ij}(start)$ to $C_{ij}(start+q-1)$ and smaller than all elements from $C_{ij}(start+q+1)$ to $C_{ij}(end)$. Therefore, we can select a single breakpoint from $C_{ij}$ by running QuickSelect once. Since we set $N_{r}=256$, the divide-and-conquer strategy can be perfectly applied to our algorithm. Specifically, a total of $\log_{2}N_{r}$ rounds need to run, and the $z$-th round select $2^{z-1}$ breakpoints by running QuickSelect in $2^{z-1}$ sub-regions generated from the (zz-1)-th round, where $z=1,...,\log_{2}N_{r}$. For each $C_{ij}$, we select the minimum element as the first breakpoint $B_{ij}(1)$ and the maximum element as the last breakpoint $B_{ij}(N_{r}+1)$ (lines 9-11). In practice, Algorithm 1 achieves 3x speedup in running time over the complete sorting scheme, as shown in Section 6.2. After getting the breakpoint set $B$, Algorithm 2 will encode all points into iSAX representations and return the set of encoded points $EP$ (lines 3-8).
Input: Parameters $K$, $L$, $n$, encoded points set $EP$, maximum size of a leaf node max_sizemax\_size
@@ 6    $target\_leaf\leftarrow$ leaf node of $T_{i}$ to insert $\langle ep_{i}(o_{z}),pos_{z}\rangle$;
8         SplitNode(target_leaftarget\_leaf);
@@ 9       $target\_leaf\leftarrow$ the new leaf node to insert $\langle ep_{i}(o_{z}),pos_{z}\rangle$;
11     Insert $\langle ep_{i}(o_{z}),pos_{z}\rangle$ to target_leaftarget\_leaf;
@@ 3 $node\leftarrow$ the $i$-th child of root node in $T$;
4 call TraverseSubtree(nodenode, $q^{\prime}$, $r^{\prime}$, $S$);
@@
Input: A node nodenode, the projected query point $q^{\prime}$, the search radius $r^{\prime}$, the set of points $S$
@@
5else if nodenode is a leaf then
6 $upper\_bound\_dist\leftarrow$ calculate the upper bound distance between $q^{\prime}$ and nodenode;
@@
8    $S\leftarrow S\,\cup$ all points in nodenode;
@@
11    while nodenode has next point do
@@ 15          $S\leftarrow S\,\cup o$;
In a DE-Tree, except the root node that has $2^{K}$ children, other internal nodes have only two children. This is because when an internal node needs to be split, we only select one of its $K$ dimensions for further bit-wise binary division. For example, in Figure 1(b), we choose the first dimension of node $^{*},00^{*}$] to split, and the representations of its two children are $0,00^{*}$] and $1,00^{*}$]. The choice of which dimension to divide is important for splitting nodes. Intuitively, splitting a node works better if the obtained two children contain similar numbers of points. Therefore, when splitting nodes, we choose the dimension that most evenly divides the points.
@@ 5 $S\leftarrow S\cup S_{i}$;
10if $\lvert\left\{o\mid o\in S\land\left\|o,q\right\|\leq c\cdot r\right\}\rvert\geq 1$ then
@@ 6    $S\leftarrow S\cup S_{i}$;
11 if $\lvert\left\{o\mid o\in S\land\left\|o,q\right\|\leq c\cdot r\right\}\rvert\geq k$ then
@@ 5.1. Quality Guarantee
Let $\mathcal{H}_{i}(o)=[h_{i1}(o),...,h_{iK}(o)]$ denote a data point $o$ in the $i$-th projected space, where i=1,,Li=1,...,L. We define three events as follows:
  • E1: If there exists a point $o$ satisfying $\left\|o,q\right\|\leq r$, then its projected distance to $q$, i.e., $\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$, is smaller than $\epsilon r$ for some i=1,,Li=1,...,L;
  • E2: If there exists a point $o$ satisfying $\left\|o,q\right\|>cr$, then its projected distance to $q$, i.e., $\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$, is smaller than $\epsilon r$ for some i=1,,Li=1,...,L;
@@
Given a point $o$ satisfying $\left\|o,q\right\|\leq r$, let $s=\left\|o,q\right\|$ and $s_{i}^{\prime}=\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$ denote the distances between $o$ and $q$ in the original space and in the $i$-th projected space, where i=1,,Li=1,...,L. From Equation 3, we have $\sqrt{\chi^{2}_{\alpha_{1}}(K)}=\epsilon$. For each independent projected space, from Lemma 6, we have $\Pr{[s_{i}^{\prime}>s\sqrt{\chi^{2}_{\alpha_{1}}(K)}]}=\Pr{[s_{i}^{\prime}>\epsilon s]}=\alpha_{1}$. Since $s\leq r$, $\Pr{[s_{i}^{\prime}>\epsilon r]}\leq\alpha_{1}$. Considering $L$ projected spaces, we have $\Pr{[\textbf{E1}]}\geq 1-\alpha_{1}^{L}=1-\frac{1}{\mathrm{e}}$. Likewise, given a point $o$ satisfying $\left\|o,q\right\|>cr$, let $s=\left\|o,q\right\|$ and $s_{i}^{\prime}=\left\|\mathcal{H}_{i}(o),\mathcal{H}_{i}(q)\right\|$ denote the distances between $o$ and $q$ in the original space and in the $i$-th projected space, where i=1,,Li=1,...,L. From Equation 3, we have $\sqrt{\chi^{2}_{\alpha_{2}}(K)}=\frac{\epsilon}{c}$. For each independent projected space, from Lemma 6, we have $\Pr{[s_{i}^{\prime}>s\sqrt{\chi^{2}_{\alpha_{2}}(K)}]}=\Pr{[s_{i}^{\prime}>\frac{\epsilon s}{c}]}=\alpha_{2}$. Since $s>cr$, i.e., $\frac{s}{c}>r$, $\Pr{[s_{i}^{\prime}>\epsilon r]}>\alpha_{2}$. Considering $L$ projected spaces, we have $\Pr{[\textbf{E2}]}\leq 1-\alpha_{2}^{L}$, thus the expected number of such points in dataset $\mathcal{D}$ is upper bounded by $(1-\alpha_{2}^{L})\cdot n$. By Markov’s inequality, we have $\Pr{[\textbf{E3}]}>1-\frac{(1-\alpha_{2}^{L})\cdot n}{\beta n}=\frac{1}{2}.$
@@
We show that when E1 and E3 hold at the same time, Algorithm 7 returns a correct $c^{2}$-$k$-ANN result. Let $o_{i}^{*}$ be the $i$-th exact nearest point to $q$ in $\mathcal{D}$, we assume that $r_{i}^{*}=\left\|o_{i}^{*},q\right\|>r_{min}$, where $r_{min}$ is the initial search radius and i=1,,ki=1,...,k. We denote the number of points in the candidate set under search radius $r$ as $\lvert S_{r}\rvert$. Obviously, when enlarging the search radius $r=r_{min},r_{min}\cdot c,r_{min}\cdot c^{2},$..., there must exist a radius $r_{0}$ satisfying $\lvert S_{r_{0}}\rvert<\beta n+k$ and $\lvert S_{c\cdot r_{0}}\rvert\geq\beta n+k$. The distribution of $r_{i}^{*}$ has three cases:
@@
    Case 1: If for all i=1,,ki=1,...,k satisfying $r_{i}^{*}\leq r_{0}$, which indicates the range queries in all $L$ index trees have been executed at $r=r_{0}$ (lines 3-8). Due to E1, all $r_{i}^{*}$ must in $S_{r_{0}}$. Since $S_{r_{0}}\subsetneqq S_{c\cdot r_{0}}$, all $r_{i}^{*}$ also must in $S_{c\cdot r_{0}}$. Therefore, Algorithm 7 returns the exact $k$ nearest points $o_{i}^{*}$ to $q$.
@@
    Case 2: If for all i=1,,ki=1,...,k satisfying $r_{i}^{*}>r_{0}$, all $r_{i}^{*}$ not belong to $S_{r_{0}}$. Since Algorithm 7 may terminate after executing range queries in part of $L$ index trees at $r=c\cdot r_{0}$ (line 8), we cannot guarantee that $r_{i}^{*}\leq c\cdot r_{0}$. However, due to E3, there are at least $k$ points $o_{i}$ in $S_{c\cdot r_{0}}$ satisfying $\left\|o_{i},q\right\|\leq c^{2}r_{0}$, i=1,,ki=1,...,k. Therefore, we have $\left\|o_{i},q\right\|\leq c^{2}r_{0}\leq c^{2}r_{i}^{*}$, i.e., each $o_{i}$ is a $c^{2}$-ANN point for corresponding $o_{i}^{*}$.
@@
    Case 3: If there exists an integer $m\in(1,k)$ such that for all i=1,,mi=1,...,m satisfying $r_{i}^{*}\leq r_{0}$ and for all i=m+1,,ki=m+1,...,k satisfying $r_{i}^{*}>r_{0}$, indicating that Case 3 is a combination of Case 1 and Case 2. For each $i\in[1,m]$, Algorithm 7 returns the exact nearest point $o_{i}^{*}$ to $q$ based on Case 1. For each $i\in[m+1,k]$, Algorithm 7 returns a $c^{2}$-ANN point for $o_{i}^{*}$ based on Case 2.